A string 4 meters along weighs 4 gm and a is stretched with a tension of 40 gm . What is the velocity of a transverse wave string?


1
Expert's answer
2021-03-29T16:47:25-0400
"v=\sqrt{\dfrac{T}{\mu}}=\sqrt{\dfrac{40\ N}{\dfrac{4\cdot10^{-3}\ kg}{4\ m}}}=200\ \dfrac{m}{s}."